OpenStudy (anonymous):

How do you get the derivative of y=3(x^2 + 2x)^2 using the chain rule. Please show step by step.

OpenStudy (zale101):

\(\large y=3(x^2 + 2x)^2 \) You first take the derivative of the outer, then take the derivative of the inside. Hint: The outer needs just the power rule and the inner needs the power rule as well. I will start by taking the derivative of the outer. \(\large \frac{dy}{dx}=3*2((x^2 + 2x))^{2-1} *\frac{d}{dx}(x^2 + 2x)\)

OpenStudy (zale101):

Can you finish it from there?

OpenStudy (zale101):

Chain rule is done my taken the derivative of the outer and multiply the inner derivatives.

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (zale101):

\(\frac{dy}{dx}=\frac{d}{dx}3(x^2 + 2x)^{2}*\frac{d}{dx}(x^2 + 2x)\)

OpenStudy (zale101):

No problem!